OpenStudy (anonymous):

Calculate the number of molecules in 6.00 moles h2s?

OpenStudy (anonymous):

1mol has avocadros number of atoms = 6.022* 10 to the power 23 6.00* 6.022*10 to the power 23=3.6132 *10 to the power 24 molecules